OpenStudy (anonymous):

On the graph of the equation 6x - y = -1, what is the value of the y-intercept?

OpenStudy (anonymous):

6x-y=-1 -y=-6x-1 (I subtracted the 6x from both sides) y=6x+1 (I divided by a negative 1) y-int is 1.